# A Weyl equidistribution theorem over function fields

### v1

@@ 1. Introduction
The quotient space $\mathcal{R}/\mathcal{Z}$ is a compact group, which we denote by $\mathcal{T}$. In analogy with the well-known isomorphism $\widehat{\mathbb{Z}}\cong\mathbb{T}=\mathbb{R}/\mathbb{Z}$, there is an isomorphism between the dual group $\widehat{\mathcal{Z}}$ and the group $\mathcal{T}$. In particular, every character on $\mathcal{Z}$ is of the form $u\mapsto e(u\alpha)$ for some $\alpha\in\mathcal{T}$, where
$e\left(\sum_{n=-\infty}^{N}c_{n}t^{n}\right)=\textup{Tr}_{\mathbb{F}_{q}/\mathbb{F}_{p}}(c_{-1}).$
@@
Moreover, $(P(u))_{u\in\mathcal{Z}}$ is well-distributed in the components of $\mathcal{F}(P)+P(K)$, $k\in K$, of $\mathcal{O}(P)$. In particular, $(P(u))_{u\in\mathcal{Z}}$ is well-distributed in $\mathcal{T}$ if and only if $\sum_{i=1}^{d}{\mathcal{F}(\eta_{i})}=\mathcal{T}$.
@@
Moreover, $(P(u_{1},\ldots,u_{d}))_{(u_{1},\ldots,u_{d})\in\mathcal{Z}^{d}}$ is well-distributed in the components of $\mathcal{F}(P)+P(K)$, $k\in K$, of $\mathcal{O}(P)$. In particular, $(P(u_{1},\ldots,u_{d}))_{(u_{1},\ldots,u_{d})\in\mathcal{Z}^{d}}$ is well-distributed in $\mathcal{T}$ if and only if $\sum_{i=1}^{m}{\mathcal{F}(\eta_{i})}=\mathcal{T}$.

### v2

@@ 1. Introduction
The quotient space $\mathcal{R}/\mathcal{Z}$ is a compact group, which we denote by $\mathcal{T}$. In analogy with the well-known isomorphism $\widehat{\mathbb{Z}}\cong\mathbb{T}=\mathbb{R}/\mathbb{Z}$, there is an isomorphism between the dual group $\widehat{\mathcal{Z}}$ and the group $\mathcal{T}$, which we now describe. Recall that, viewing $\mathbb{F}_{q}$ as an extension of $\mathbb{F}_{p}$, there is a trace map $\textup{Tr}_{\mathbb{F}_{q}/\mathbb{F}_{p}}:\mathbb{F}_{q}\to\mathbb{F}_{p}$, which can be written explicitly as $\textup{Tr}_{\mathbb{F}_{q}/\mathbb{F}_{p}}(c)=c+c^{p}+\ldots+c^{p^{k-1}}$, where $q=p^{k}$. Every character on $\mathcal{Z}$ is then of the form $u\mapsto e(u\alpha)$ for some $\alpha\in\mathcal{T}$, where
$e\left(\sum_{n=-\infty}^{N}c_{n}t^{n}\right)=\exp\left(\frac{2\pi i\cdot\textup{Tr}_{\mathbb{F}_{q}/\mathbb{F}_{p}}(c_{-1})}{p}\right).$
@@
Remark 1.3.
A proof of Theorem 1.2 by different methods appears in [CGLLW25].
@@
Moreover, $(P(u))_{u\in\mathcal{Z}}$ is well-distributed in the components of $\mathcal{F}(P)+P(k)$, $k\in K$, of $\mathcal{O}(P)$. In particular, $(P(u))_{u\in\mathcal{Z}}$ is well-distributed in $\mathcal{T}$ if $\sum_{i=1}^{d}{\mathcal{F}(\eta_{i})}=\mathcal{T}$.
Remark 2.2.
There is a subtle difference between the formulation of [BL16, Theorem 0.3] as given in Theorem 2.1 and the original formulation from [BL16]. In the present paper, we write that $\sum_{i=1}^{d}{\mathcal{F}(\eta_{i})}=\mathcal{T}$ is a sufficient condition for $(P(u))_{u\in\mathcal{Z}}$ to be well-distributed in $\mathcal{T}$. In [BL16], it is incorrectly claimed that this condition is also necessary, a claim that was duplicated in an earlier version of this article on arXiv. (The word “if” in the final sentence of Theorem 2.1 was previously the phrase “if and only if” and appears in [BL16] as “iff.” The “if” direction follows from the prior part of the theorem statement, while the “only if” direction does not.) A consequence of the incorrect statement is distilled as Hypothesis 5.1 in [CGLLW25], where a counterexample is also produced. Since the present paper is concerned only with a sufficient condition for equidistribution, this small error does not impact our usage of Theorem 2.1. Similar comments apply to the multivariable equidistribution result in Theorem 3.2.
@@
Moreover, $(P(u_{1},\ldots,u_{d}))_{(u_{1},\ldots,u_{d})\in\mathcal{Z}^{d}}$ is well-distributed in the components of $\mathcal{F}(P)+P(k)$, $k\in K$, of $\mathcal{O}(P)$. In particular, $(P(u_{1},\ldots,u_{d}))_{(u_{1},\ldots,u_{d})\in\mathcal{Z}^{d}}$ is well-distributed in $\mathcal{T}$ if $\sum_{i=1}^{m}{\mathcal{F}(\eta_{i})}=\mathcal{T}$.